OpenStudy (anonymous):

tan^-1(2x) + tan^-1(3x) = π/4 find x... pls help

OpenStudy (anonymous):

\[tan^{-1} \frac{5x}{1 - 6x^2} = \frac{\pi}{4}\]

OpenStudy (anonymous):

should we use tan(A+B) ?

OpenStudy (anonymous):

1 - 6x^2 = 5x 6x^2 + 5x -1 =0 (6x - 1)(x + 1) =0 x = -1 , 1/6

OpenStudy (anonymous):

yea we should

OpenStudy (anonymous):

thanks a lot!! :) can u pls tell how u got 1-6x^2 =5x

OpenStudy (anonymous):

\[tan^{-1} a + tan^{-1} b = tan^{-1} \frac{ a + b}{1 - ab}\] the real thing is down here ...i wrote the above by mistake \[tan^{-1} x = y\] \[x = tany\]

OpenStudy (anonymous):

okie...tnx.. :)